OpenStudy (anonymous):

I GIVE MEDALS. A car drives on a circular track with a constant speed of 84 m/s. The track has a radius of 225 m. What is the centripetal acceleration of the car?

OpenStudy (anonymous):

f = mv^2/r ma = mv^2/r a = v^2/r a = 84^2/225 a = 31.36 m/s^2

OpenStudy (anonymous):

How would i find the acceleration of the car measured in g? @li_za12

OpenStudy (cybershadow):

what do u mean Exactly by "acceleration of the car measured in g". do u want to show this acceleartion in terms of g? :S if that so then--- a = v^2/r a = 31.36 m/s^2 also we know g = 10 m/s^2 a/g = 31.36/10 a = 31.36/10 g m/s^2 a =3.13g m/s^2 Was this ur question? :S

OpenStudy (anonymous):

yes, thanks! How would i figure out what speed would give the car an acceleration of 9.8 m/s2? @CyberShadow

OpenStudy (cybershadow):

oh that , here it is --- As we know that , a = v^2/r and we want to get an acceleration of a = 9.8 m/s^2 also r = 225 m therefore 9.8 = v^2/225 implies v^2 = 9.8 * 225 v = (2205)^1/2 v = 46.9 m/s get it? :)

OpenStudy (anonymous):

That makes so much more sense, haha. thanks! @CyberShadow

OpenStudy (cybershadow):

No worries :)